OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 1024 and a3 = 64?

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf a_{1} \times r=a_{2} }\) and \(\large\color{blue}{ \bf a_{3} \div r=a_{2} }\) knowing that they are both equal to a2, (and knowing a1 and a3) can you find the r ?

OpenStudy (anonymous):

1,024 x r = a_2 64/r = a_2 So... my guess is to divide 1,024 by 64 So the answer is 0.0625

OpenStudy (solomonzelman):

We can set the equations as, \(\large\color{blue}{ \bf a_1 \times r = a_3 \div r }\) because each side gives the value of the second term.

OpenStudy (anonymous):

Thank you again :) I just need to be shown the formula and see what matches to what and I think I'm set! :) Could you help me with one last question?

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf 2014 \times r = 64 \div r }\) \(\large\color{blue}{ \bf 2014 \times r^2 = 64 }\) \(\large\color{blue}{ \bf r^2 = \frac{64}{2014} }\)

OpenStudy (solomonzelman):

So the ratio is the square root of that fraction... I don't like dealing with not-perfect squares -:(

OpenStudy (anonymous):

oh! So that wasn't the correct answer?

OpenStudy (solomonzelman):

https://www.google.com/#q=64+divided+by+2014 and then square root of that, or you can work with the exact value. Whatever you prefer.

OpenStudy (anonymous):

ok thank you :) you're a lifesaver!

OpenStudy (solomonzelman):

Lets work with an exact value. But before can you answer one question ?

OpenStudy (solomonzelman):

Do you understand why (ready? pay attention!!! ) are there going to be different ratios, and why don't we have to worry about this (and we can just work with the positive ratio) ?

OpenStudy (solomonzelman):

2 different ratios...